# The Benefits of a Concise Chain of Thought on Problem-Solving in Large Language Models

### v1

@@ Abstract
In this paper, we introduce Concise Chain-of-Thought (CCoT) prompting. We compared standard CoT and CCoT prompts to see how conciseness impacts response length and correct-answer accuracy. We evaluated this using GPT-3.5 and GPT-4 with a multiple-choice question-and-answer (MCQA) benchmark. CCoT reduced average response length by 48.70% for both GPT-3.5 and GPT-4 while having a negligible impact on problem-solving performance. However, on math problems, GPT-3.5 with CCoT incurs a performance penalty of 27.69%. Overall, CCoT leads to an average per-token cost reduction of 22.67%. These results have practical implications for AI systems engineers using LLMs to solve real-world problems with CoT prompt-engineering techniques. In addition, these results provide more general insight for AI researchers studying the emergent behavior of step-by-step reasoning in LLMs. All code, data, and supplemental materials are available on GitHub [^1] https://github.com/matthewrenze/jhu-concise-cot.
1 Introduction
1.1 Background
@@
However, to create effective LLM solutions, prompt engineering is often necessary. As a result, AI systems engineers have developed various techniques to improve the performance of LLMs for specific use cases and problem domains. These techniques include Chain of Thought (CoT) prompting Mialon et al. (2023); White et al. (2023).
1.2 Chain-of-Thought (CoT) Prompting
CoT is a prompt engineering technique that instructs the LLM to reason through a problem in a step-by-step manner. Reasoning step-by-step increases the likelihood of the LLM producing a correct solution to the problem. As a result, CoT improves LLM performance on many problem-solving tasks Kojima et al. (2022); Wei et al. (2022); Zhou et al. (2023).
There are multiple versions of CoT prompting with various pros and cons. Zero-shot CoT instructs the LLM to “think step-by-step” through the problem in the system prompt Kojima et al. (2022); Zhou et al. (2023). Few-shot CoT provides a series of examples as problem-solution pairs with the CoT explicitly stated in each example solution Wei et al. (2022).
CoT prompting has been shown to improve LLM performance by up to 80% for certain problem tasks and problem domains Kojima et al. (2022). However, this performance increase comes at the expense of increased LLM response length. As a result, the cost of using the LLM with CoT grows in proportion to response length.
1.3 Concise Prompting
@@
There are two main implementations of concise prompting. Zero-shot prompting instructs the LLM to “be concise” in its response Crispino et al. (2023); Kadous (2023). Few-shot prompting requires the prompt engineer to create a series of problem-solution example pairs with concisely written text in each example solution.
While concise prompting is beneficial for reducing resource costs, it may negatively impact the performance of the LLM on some problem-solving tasks Crispino et al. (2023). This is because the LLM requires additional verbosity to fully elaborate the steps in its thought process to produce a correct solution.
@@
1.4 Concise Chain-of-Thought (CCoT)
@@
CCoT is achieved by instructing the LLM to both ”think step-by-step” and ”be concise”. In addition, the LLM is provided with few-shot examples that include a sample problem and a concise solution rather than a standard (i.e., verbose) solution.
For an example of the CCoT prompt and few-shot examples, please see Appendix B.3 of the supplemental materials.
1.5 Hypotheses
To test the impact of CCoT on an LLM’s response length, we developed the following hypotheses:
  • Response-Length Null Hypothesis (RL-H0): CCoT has no effect on the number of response tokens produced by the LLM compared to standard CoT prompting.
  • Response-Length Alternative Hypothesis (RL-H1): CCoT decreases the number of response tokens produced by the LLM compared to standard CoT prompting.
To test the impact of CCoT on an LLM’s problem-solving performance, we developed the following hypotheses:
  • Performance Null Hypothesis (P-H0): CCoT has no effect on the correct-answer accuracy of the LLM compared to standard CoT prompting.
  • Performance Alternative Hypothesis (P-H1): CCoT decreases the correct-answer performance of the LLM compared to standard CoT prompting.
To determine the validity of these hypotheses, we performed a hypothesis test using a significance level $(\alpha)$ of 0.05.
1.6 Significance
These results have practical implications for AI systems engineers using LLMs with CoT in their solutions.
If CCoT reduces response length (i.e., if RL-H1 is true), then AI systems engineers can use CCoT to reduce LLM costs. Third-party LLM APIs are typically priced per token Anyscale (2023); OpenAI (2023e). So, reducing response length will reduce total costs. Reducing response length also reduces energy consumption and response wait times.
In addition, if CCoT does not decrease performance (i.e., if P-H0 is true), then there is no performance penalty for implementing CCoT. As a result, AI systems engineers should prefer CCoT over standard CoT.
These results also have theoretical implications for AI researchers studying CoT reasoning in LLMs.
If we can reduce the length of a CoT without impacting performance, then only some aspects of a CoT are relevant to the LLM’s problem-solving performance. This discovery raises new questions about which specific tokens or aspects of an LLM’s CoT are necessary vs. which are superfluous.
1.7 Prior Literature
Few-shot CoT prompting was introduced by Wei et al. in Jan 2022 Wei et al. (2022). Zero-shot CoT was introduced four months later, in May 2022, by Kojima et al. Kojima et al. (2022). The zero-shot CoT method was then further refined by Zhou et al. using the Automatic Prompt Engineer (APE) method in Oct 2022 Zhou et al. (2023).
In Sept 2022, Madaan and Yazdanbakhsh attempted to separate symbols, patterns, and text to determine their individual effects on CoT and develop their own concise CoT prompt technique Madaan and Yazdanbakhsh (2022a). However, as of Jan 2023, their paper was withdrawn from publication due to technical issues Madaan and Yazdanbakhsh (2022b).
@@ 2.1 Data
We reviewed existing literature to identify a set of candidate benchmarks that spanned multiple problem domains and difficulty levels. The source benchmarks we selected consist of ARC, AGIEval, HellaSwag, and MedMCQA Clark et al. (2018); Zellers et al. (2019); Pal et al. (2022); Zhong et al. (2023); Liu et al. (2020); Wang et al. (2021).
@@
For a complete list of the source problem sets used to create the MCQA test set, see Appendix A.1 in the supporting materials. For a sample MCQA problem, see Appendix A.2.
2.2 Models
We used two popular LLMs to test our hypotheses – namely GPT-3.5 and GPT-4.
GPT-3.5 is a Generative Pre-trained Transformer (GPT) created by OpenAI and publicly released as “ChatGPT” in Nov 2022 Brown et al. (2020); OpenAI (2022). GPT-4 is a more powerful GPT with more advanced capabilities OpenAI (2023b, a). However, GPT-4 costs roughly 30x more per output token than GPT-3.5 OpenAI (2023e).
GPT-3.5 and GPT-4 can be accessed via an Application Programming Interface (API) hosted by OpenAI or Microsoft. We used Microsoft Azure Open AI Service for our experiments – though our results should hold for either platform – since they both use the same underlying foundational models Microsoft (2023); OpenAI (2023c, d).
2.3 Prompts
We used three prompt-engineering techniques to test our hypotheses. These prompts consisted of an answer-only prompt, a standard (i.e., verbose) CoT prompt, and a concise CoT prompt.
@@ 2.3 Prompts
For samples of all three prompts, see Appendix B in the supplemental materials.
2.4 Metrics
To test our hypotheses, we measured the LLM’s response length and correct-answer accuracy.
Response length was measured as the number of output tokens produced by the LLM in its response. A token is the smallest unit of information an LLM can process. Both GPT-3.5 and GPT-4 work with sub-word tokens. On average, there are approximately 1.3 tokens per word OpenAI (2023f).
@@ 2.4 Metrics
2.5 Analysis
To test the response-length hypothesis (RL-H0), we performed a Mann-Whitney U (MWU) test on the means of the CoT and CCoT response lengths. To test our performance hypothesis (P-H0), we performed an MWU test on the means of the correct-answer accuracy for standard CoT and CCoT Mann and Whitney (1947); The SciPy Community (2023).
@@ 3 Results
Figure 3: GPT-3.5 with CCoT decreases response length significantly across all problem domains compared to CoT.
Figure 4: GPT-4 with CCoT decreases response length significantly across all problem domains compared to CoT.
3.1 Response Length
@@
Given a significance level of 0.05, the low p-values indicate that the differences in response lengths between CCoT and CoT are statistically significant for both GPT-3.5 and GPT-4. Thus, we have sufficient evidence to reject the null hypothesis (RL-H0) in favor of the alternative hypothesis (RL-H1).
@@
A visual analysis of response length by exam shows this decrease in response length holds across all problem domains. This pattern can be observed for both GPT-3.5 and GPT-4 (see Figures 3 and 4).
3.2 Performance
For GPT-3.5, we saw a 2.95% decrease in average accuracy when comparing CCoT to CoT. The MWU test yielded $(U=507,648.50,p=0.55)$, indicating the differences in the means were not statistically significant.
For GPT-4, we saw an increase in response length of 0.25% when comparing CCoT to CoT prompting. The MWU test yielded $(U=485,396.00,p=0.21)$, also indicating a non-significant difference in the means.
Given our significance level of 0.05, the high p-values indicate that the differences in accuracy between CCoT and CoT are not statistically significant for either GPT-3.5 or GPT-4. Thus, we do not have sufficient evidence to reject the null hypothesis for performance (P-H0).
A visual analysis of these data supports these findings (see Figures 6 and 6)
A more in-depth analysis by exam revealed that GPT-3.5 with CCoT produced a statistically significant decrease in performance only for math problems (i.e., the aqua-rat-100 and sat-math-100 exams) compared to CoT. Other problem domains saw no decrease in accuracy (see Figure 7).
For these two math exams, GPT-3.5 with CCoT resulted in an average reduction in accuracy of 27.69%. An MWU test on the means of CoT and CCoT yielded $(U=26,546.00,p<0.001)$. Thus, for the special case of math problems, we must reject the null hypothesis (P-H0) in favor of the alternative hypothesis (P-H1).
On the other hand, GPT-4 with CCoT did not result in a statistically significant decrease in performance on math problems compared to CoT. Both CCoT and CoT had roughly equivalent performance on all problem domains. CCoT also significantly outperformed the answer-only prompt on math problems (see Figure 8).
An MWU test on the means yielded $(U=21,471.50,p=0.17)$. Thus, in the special case of the mathematics problem domain, the null hypothesis (P-H0) still holds for GPT-4.
@@ 3.2 Performance
Figure 7: Verbose CoT prompting with GPT-3.5 significantly improved performance on math problems (i.e., aqua-rat-100 and sat-math-100) compared to CCoT but had minimal impact on other problem domains.
Figure 8: Verbose CoT and CCoT prompting with GPT-4 significantly improved accuracy for math problems (i.e., aqua-rat-100 and sat-math-100) but had minimal impact on other problem domains.
3.3 Cost Analysis
GPT-3.5 with CoT (\$) GPT-3.5 with CCoT (\$) GPT-4 with CoT (\$) GPT-4 with CCoT (\$)
Input Cost 0.55 0.51 16.37 15.29
Output Cost 0.33 0.17 10.53 5.29
Total Cost 0.88 0.69 26.90 20.58
Cost Savings 21.85% 23.49%
Table 1: CCoT significantly reduces total costs per 1,000 problems for both GPT-3.5 and GPT-4.
To understand the practical implications of these results, we computed the cost of solving all 1,000 problems using the current per-token pricing model for GPT-3.5 and GPT-4222For our study, we solved each MCMQ problem 10 times. So, our total cost was roughly 10x the values we are reporting..
Currently, GPT-3.5 is priced at \$0.001 per 1,000 input tokens and \$0.002 per 1K output tokens [^3] All prices are in US Dollars (USD).. GPT-4 is priced at \$0.03 per 1K input tokens and \$0.06 per 1K output tokens OpenAI (2023e). As a result, CCoT produced a total cost savings of 21.85% for GPT-3.5 and 23.49% for GPT-4 (see Table 1). These cost savings should scale linearly.
4 Discussion
4.1 Limitations
@@
First, our study involved only two LLMs – both of which were versions of GPT. As a result, these findings may not be replicable using other proprietary and open-source LLMs like Llama 2, PaLM, and Claude.
@@
4.2 Implications
These results have practical implications for AI systems engineers building AI systems with LLMs.
Since most proprietary LLM APIs charge on a per-token pricing model, reducing the number of output tokens in an LLM’s response has direct cost savings Anyscale (2023); OpenAI (2023e). In addition, getting more concise responses without sacrificing accuracy has direct savings in energy consumption and response wait times.
@@ 4.2 Implications
4.3 Future Research
To improve upon this research, we suggest the following follow-up experiments:
@@ 4.3 Future Research
Second, we recommend testing more CCoT prompt variations. Instructing the LLM to be even more concise in the system prompt may further decrease response length. In addition, even more concisely written few-shot examples may produce even more succinct responses without sacrificing accuracy.
@@ 5 Conclusion
We demonstrated that it reduced response token length for GPTs by 48.70% while performing as well as standard CoT. GPT-4 incurred no performance penalty in any problem domain. However, for math problems, GPT-3.5 incurred a 27.69% reduction in accuracy.
@@
These results have practical implications for AI systems engineers using CoT to build problem-solving AI systems. They also have theoretical implications for AI researchers studying LLM reasoning processes.

### v2

@@ Abstract
In this paper, we introduce Concise Chain-of-Thought (CCoT) prompting. We compared standard CoT and CCoT prompts to see how conciseness impacts response length and correct-answer accuracy. We evaluated this using GPT-3.5 and GPT-4 with a multiple-choice question-and-answer (MCQA) benchmark. CCoT reduced average response length by 48.70% for both GPT-3.5 and GPT-4 while having a negligible impact on problem-solving performance. However, on math problems, GPT-3.5 with CCoT incurs a performance penalty of 27.69%. Overall, CCoT leads to an average per-token cost reduction of 22.67%. All code, data, and supplemental materials are available on GitHub at https://github.com/matthewrenze/jhu-concise-cot
Index Terms:
large language model, LLM, chain-of-thought, CoT, concise
I Introduction
I-A Background
@@
However, to create effective LLM solutions, prompt engineering is often necessary. So, AI systems engineers have developed various techniques to improve the performance of LLMs for specific use cases and problem domains. These techniques include Chain of Thought (CoT) prompting [1, 2].
I-B Chain-of-Thought (CoT) Prompting
CoT is a prompt engineering technique that instructs the LLM to reason through a problem in a step-by-step manner. Reasoning step-by-step increases the likelihood of the LLM producing a correct solution to the problem. As a result, CoT improves LLM performance on many problem-solving tasks [3, 4, 5].
There are multiple versions of CoT prompting with various pros and cons. Zero-shot CoT instructs the LLM to “think step-by-step” through the problem in the system prompt [3, 5]. Few-shot CoT provides a series of examples as problem-solution pairs with the CoT explicitly stated in each example solution [4].
CoT prompting has been shown to improve LLM performance by up to 80% for certain problem tasks and problem domains [3]. However, this performance increase comes at the expense of increased LLM response length. As a result, the cost of using the LLM with CoT grows in proportion to response length.
I-C Concise Prompting
@@
There are two main implementations of concise prompting. Zero-shot prompting instructs the LLM to “be concise” in its response [6, 7]. Few-shot prompting requires the prompt engineer to create a series of problem-solution example pairs with concisely written text in each example solution.
While concise prompting is beneficial for reducing resource costs, it may negatively impact the performance of the LLM on some problem-solving tasks [6]. This is because the LLM requires additional verbosity to fully elaborate the steps in its thought process to produce a correct solution.
@@
I-D Concise Chain-of-Thought (CCoT)
@@
CCoT is achieved by instructing the LLM to both "think step-by-step" and "be concise". In addition, the LLM is provided with few-shot examples that include a sample problem and a concise solution rather than a standard (i.e., verbose) solution.
For an example of the CCoT prompt and few-shot examples, please see Figure 12 in the appendix.
I-E Contribution
Our research builds upon prior literature by determining how concise prompting impacts CoT prompting. These results have practical implications for AI engineers building software solutions with LLMs.
First, if CCoT reduces response length, then AI systems engineers can use CCoT to reduce LLM costs. Third-party LLM APIs are typically priced per token [8, 9]. So, reducing response length will reduce total costs. Reducing response length also reduces energy consumption and response wait times.
Second, if CCoT does not decrease performance, then there is no performance penalty for implementing CCoT. As a result, AI systems engineers should prefer CCoT over standard CoT.
Finally, these results have theoretical implications for AI researchers studying CoT reasoning in LLMs. If we can reduce the length of a CoT without impacting performance, then only some aspects of a CoT are relevant to the LLM’s problem-solving performance. This discovery raises new questions about which specific tokens or aspects of an LLM’s CoT are necessary vs. which are superfluous.
I-F Prior Literature
Few-shot CoT prompting was introduced by Wei et al. in Jan 2022 [4]. Zero-shot CoT was introduced four months later, in May 2022, by Kojima et al. [3]. The zero-shot CoT method was then further refined by Zhou et al. using the Automatic Prompt Engineer (APE) method in Oct 2022 [5].
In Sept 2022, Madaan and Yazdanbakhsh attempted to separate symbols, patterns, and text to determine their individual effects on CoT and develop their own concise CoT prompt technique [10]. However, as of Jan 2023, their paper was withdrawn from publication due to technical issues [11].
@@ II-A Data
We reviewed existing literature to identify a set of candidate benchmarks that spanned multiple problem domains and difficulty levels. The source benchmarks we selected consist of ARC, AGIEval, HellaSwag, and MedMCQA [12, 13, 14, 15, 16, 17].
@@
For a complete list of the source problem sets used to create the MCQA test set, see Table II in the appendix. For a sample MCQA problem, see Figure 9 in the appendix.
II-B Models
We used two popular LLMs in our experiment – namely GPT-3.5 and GPT-4.
GPT-3.5 is a Generative Pre-trained Transformer (GPT) created by OpenAI and publicly released as “ChatGPT” in Nov 2022 [18, 19]. GPT-4 is a more powerful GPT with more advanced capabilities [20, 21]. However, GPT-4 costs roughly 30x more per output token than GPT-3.5 [9].
GPT-3.5 and GPT-4 can be accessed via an Application Programming Interface (API) hosted by OpenAI or Microsoft. We used Microsoft Azure OpenAI Service for our experiments – though our results should hold for either platform – since they both use the same underlying foundational models [22, 23, 24].
II-C Prompts
We used three prompt-engineering techniques in our experiment. These prompts consisted of an answer-only prompt, a standard (i.e., verbose) CoT prompt, and a concise CoT prompt.
@@ II-C Prompts
For samples of all three prompts, see Figures 10, 11, and 12 in the appendix.
II-D Metrics
To compare the performance of our prompts, we measured the LLM’s response length and correct-answer accuracy.
Response length was measured as the number of output tokens produced by the LLM in its response. A token is the smallest unit of information an LLM can process. Both GPT-3.5 and GPT-4 work with sub-word tokens. On average, there are approximately 1.3 tokens per word [25].
@@ II-D Metrics
II-E Analysis
To determine the statistical significance of our results, we performed a set of Mann-Whitney U (MWU) tests [26, 27].
For the response-length results, we performed an MWU test on the means of the CoT and CCoT response lengths. For the performance results, we performed an MWU test on the means of the correct-answer accuracy for standard CoT and CCoT.
@@ III Results
III-A Response Length
@@
A further analysis of response length by exam shows this decrease in response length holds across all problem domains. This pattern can be observed for both GPT-3.5 and GPT-4 (see Figures 5 and 6 in the appendix).
III-B Performance
@@
For GPT-3.5, we saw a 2.95% decrease in average accuracy when comparing CCoT to CoT. The MWU test yielded $(U=507,648.50,p=0.55)$, indicating the difference in the means was not statistically significant.
For GPT-4, we saw an increase in average accuracy of 0.25% when comparing CCoT to CoT prompting. The MWU test yielded $(U=485,396.00,p=0.21)$, also indicating a non-significant difference in the means.
A visual analysis of these data supports these findings (see Figures 4 and 4)
A more in-depth analysis accuracy by exam revealed that GPT-3.5 with CCoT produced a statistically significant decrease in performance only for math problems (i.e., AQUA-RAT and SAT Math) compared to CoT. Other problem domains saw no decrease in accuracy (see Figure 7 in the appendix).
For these two math exams, GPT-3.5 with CCoT resulted in an average reduction in accuracy of 27.69%. An MWU test on the means of CoT and CCoT yielded $(U=26,546.00,p<0.001)$. Thus, for GPT-3.5, only, math problems appear to be a special case where conciseness negatively impacts performance.
On the other hand, GPT-4 with CCoT did not result in a statistically significant decrease in performance on math problems compared to CoT. Both CCoT and CoT had roughly equivalent performance on all problem domains. CCoT also significantly outperformed the answer-only prompt on math problems (see Figure 8 in the appendix).
III-C Cost Analysis
To understand the practical implications of these results, we computed the cost of solving all 1,000 problems using the current per-token pricing model for GPT-3.5 and GPT-4111For our study, we solved each MCMQ problem 10 times. So, our total cost was roughly 10x the values we are reporting..
Currently, GPT-3.5 is priced at \$0.001 per 1,000 input tokens and \$0.002 per 1K output tokens [^2] All prices are in US Dollars (USD).. GPT-4 is priced at \$0.03 per 1K input tokens and \$0.06 per 1K output tokens [9]. As a result, CCoT produced a total cost savings of 21.85% for GPT-3.5 and 23.49% for GPT-4 (see Table I). These cost savings should scale linearly.
GPT-3.5 GPT-3.5 GPT-4 GPT-4
CoT CCoT CoT CCoT
Input Cost (\$) 0.55 0.51 16.37 15.29
Output Cost (\$) 0.33 0.17 10.53 5.29
Total Cost (\$) 0.88 0.69 26.90 20.58
Cost Savings (%) 21.85 23.49
TABLE I: CCoT significantly reduces total costs per 1,000 problems for both GPT-3.5 and GPT-4.
IV Discussion
IV-A Limitations
@@
First, our study involved only two LLMs – both of which were versions of OpenAI’s GPT. As a result, these findings may not generalize to other proprietary and open-source LLMs like Llama 2, PaLM, and Claude.
@@
IV-B Implications
These results have practical implications for AI engineers building LLM-based software solutions.
Since most proprietary LLM APIs charge on a per-token pricing model, reducing the number of output tokens in an LLM’s response has direct cost savings [8, 9]. In addition, getting more concise responses without sacrificing accuracy has direct savings in energy consumption and response times.
@@ IV-B Implications
IV-C Future Research
To improve upon this research, we suggest additional follow-up experiments:
@@ IV-C Future Research
Second, we recommend testing more CCoT prompt variations. Instructing the LLM to be even more concise in the system prompt may further decrease response length. In addition, more concisely written few-shot examples may produce even more succinct responses without sacrificing accuracy.
@@ V Conclusion
We demonstrated that CCoT prompting reduced response token length for GPTs by 48.70% while performing as well as standard CoT. GPT-4 incurred no performance penalty in any problem domain. However, for math problems, GPT-3.5 incurred a 27.69% reduction in accuracy.
@@
Appendix A Prompts
A-A Answer-Only Prompt
[System Prompt]
You are an intelligent assistant.
Your task is to answer the following multiple-choice questions.
You MUST answer the question using the following format ’Action: Answer("[choice]")’
The parameter [choice] is the letter or number of the answer you want to select (e.g. "A", "B", "C", or "D").
For example, ’Answer("C")’ will select choice "C" as the best answer.
You MUST select one of the available choices; the answer CANNOT be "None of the Above".
[Example Problem]
Question: What is the capital of the state where Johns Hopkins University is located?
Choices:
  A: Baltimore
  B: Annapolis
  C: Des Moines
  D: Las Vegas
[Example Solution]
Action: Answer("B")
Figure 10: Sample of the answer-only system prompt and one-shot example
A-B Verbose CoT Prompt
[System Prompt]
You are an intelligent assistant.
Your task is to answer the following multiple-choice questions.
Think step-by-step through the problem to ensure you have the correct answer.
Then, answer the question using the following format ’Action: Answer("[choice]")’
The parameter [choice] is the letter or number of the answer you want to select (e.g. "A", "B", "C", or "D").
For example, ’Answer("C")’ will select choice "C" as the best answer.
You MUST select one of the available choices; the answer CANNOT be "None of the Above".
[Example Problem]
Question: What is the capital of the state where Johns Hopkins University is located?
Choices:
A: Baltimore
B: Annapolis
C: Des Moines
D: Las Vegas
[Example Solution]
Thought:
Johns Hopkins University is located in Baltimore.
Baltimore is a city located in the State of Maryland.
The capital of Maryland is Annapolis.
Therefore, the capital of the state where Johns Hopkins University is located is Annapolis.
The answer is B:
Figure 11: Sample of standard (i.e., verbose) CoT system prompt and one-shot example
A-C Concise CoT Prompt
[System Prompt]
You are an intelligent assistant.
Your task is to answer the following multiple-choice questions.
Think step-by-step through the problem to ensure you have the correct answer.
Then, answer the question using the following format ’Action: Answer("[choice]")’
The parameter [choice] is the letter or number of the answer you want to select (e.g. "A", "B", "C", or "D").
For example, ’Answer("C")’ will select choice "C" as the best answer.
You MUST select one of the available choices; the answer CANNOT be "None of the Above".
Be concise.
[Example Problem]
Question: What is the capital of the state where Johns Hopkins University is located?
Choices:
A: Baltimore
B: Annapolis
C: Des Moines
D: Las Vegas
[Example Solution]
Thought:
Johns Hopkins University is located in Baltimore, Maryland.
The capital of Baltimore is Annapolis.
Action: Answer("B")
Figure 12: Sample of CCoT system prompt and one-shot example